# Confidence-Aware Multi-Field Model Calibration

### v1

@@ 4.1.1. Settings
Dataset Type Method AUCnormal-$\uparrow$ LogLossnormal-$\downarrow$ Field-RCEnormal-$\downarrow$ Multi-field-RCEnormal-$\downarrow$ ECEnormal-$\downarrow$ MVCEnormal-$\downarrow$
@@
AliExpress (CTCVR) No Calib. N/A 0.8773 0.0046 0.3274 0.3273 2.085e-4 2.085e-4
Binning HistoBin 0.8752 0.0047 0.1417 0.1411 8.688e-5 8.742e-5
IsoReg 0.8770 0.0046 0.1397 0.1394 9.811e-5 8.568e-5
Scaling Naive 0.8773 0.0046 0.1431 0.1429 8.379e-5 8.450e-5
PlattScaling 0.8773 0.0045 0.1406 0.1403 8.597e-5 8.658e-5
GaussCalib 0.8773 0.0045 0.1433 0.1430 8.784e-5 8.836e-5
GammaCalib 0.8773 0.0045 0.1404 0.1401 8.538e-5 8.600e-5
Hybrid SIR 0.8773 0.0046 0.3299 0.3298 2.102e-4 2.102e-4
NeuCalib 0.8773$\pm$0.0003 0.0046$\pm$5e-7 0.1499$\pm$0.0480 0.1497$\pm$0.0480 1.068e-4$\pm$3.236e-5 8.617e-5$\pm$3.300e-5
AdaCalib 0.8759$\pm$0.0004 0.0046$\pm$5e-8 0.1459$\pm$0.0383 0.1458$\pm$0.0375 1.045e-4$\pm$3.131e-5 8.003e-5$\pm$3.316e-5
ConfCalib 0.8778 0.0045 0.1360 0.1360 8.352e-5 8.425e-5
@@
4.1.2. Datasets
We conduct experiments on our industrial dataset, as well as two public datasets, including Avazu  [^2] https://kaggle.com/competitions/avazu-ctr-prediction and AliExpress (Li et al., 2020). The data statistics are shown in Table 1.
Avazu is a popular dataset on mobile ad click-through rate (CTR) prediction with 40.4 million samples in 10 days ordered by the impression time. We split the dataset by day in a ratio of 7:1:2 for training, validation, and test. For our multi-field joint calibration method, the target fields are set to ”site_id” (3480) [^3] Represent the number of possible values of the corresponding field., ”app_id” (4864) and ”app_domain” (305). For other methods considering a single field, the target field is set to ”site_id”.
AliExpress is a dataset for multi-task learning of click-through rate and conversion rate prediction in e-commerce. Here, we choose the data from four countries except Russia and only use the conversion label to build a single click-through conversion rate (CTCVR) prediction task. The training, validation, and test set are split in a ratio of 6:1:3, totaling 103.8 million samples. The timestamp of the test set is strictly later than the other two sets. For our method, the target fields are ”country” (4) and an anonymous field, ”feature_1” (10). For other methods, the target field is ”feature_1”.
For the industrial data, we collect 38.8M clicked samples in 10 consecutive days from the log of our ad platform to build a post-click conversion rate (CVR) prediction task with a split ratio of 7:1:2 by day. For our method, the target fields for calibration are ”advertiser_id” (529), ”app_size” (82), and ”app_id” (56). For other methods, the target field is ”advertiser_id”.
The ratios of validation sets are relatively small while splitting, which is more in line with the short time window for data collection during online running.
4.1.3. Comparative Methods
We choose several representative methods of the three categories of methods for comparison. The scaling-based methods include naive scaling (multiply all predicted scores by a value $k$ to make their mean equal to the positive sample ratio), Platt Scaling (PlattScaling) (Platt et al., 1999), Gaussian Calibration (GaussCalib) (Kweon et al., 2022) and Gamma Calibration (GammaCalib) (Kweon et al., 2022). The binning-based methods include Histogram Binning (HistoBin) (Zadrozny and Elkan, 2001) and Isotonic Regression (IsoReg) (Zadrozny and Elkan, 2002). The hybrid methods include Smoothed Isotonic Regression (SIR) (Deng et al., 2021), Neural Calibration (NeuCalib) (Pan et al., 2020) and AdaCalib (Wei et al., 2022), with the latter two methods based on neural networks. The number of bins for NeuCalib and AdaCalib is set according to the initial settings when proposed.
4.1.4. Evaluation Metrics
Referring to prior methods, we continue to use field-level relative calibration error (Field-RCE) (Pan et al., 2020) to evaluate the calibration error on a certain field, which formula is
@@ 4.1.4. Evaluation Metrics
where $\mathcal{D}$ is the dataset, $\mathcal{Z}$ is the possible value set of the target field and $\mathcal{D}_{z}$ is the subset of samples with the target field value $z$ satisfying $\sum_{z\in\mathcal{Z}}{|\mathcal{D}_{z}|}=|\mathcal{D}|$. Intuitively, a lower Field-RCE indicates lower biases along the target field. Furthermore, we use the mean Field-RCE on multiple fields (Multi-field-RCE) to measure the multi-field calibration performance.
where $R$ is the number of shuffle operations, $T$ is the number of bins, and $\mathcal{D}_{r,t}$ is the subset of samples in the $t$-th bin after the $r$-th shuffle. MVCE allows us to inspect the calibration error from a more global perspective beyond the level of feature fields, thereby highlighting the effectiveness of multi-field calibration. The above calibration error metrics focus on different aspects, thus allowing us to examine the performance of each method more comprehensively.
Additionally, we also report AUC and LogLoss to evaluate the ranking performance after calibration since our method and main comparative methods are not globally isotonic.
4.2. Results
Figure 5. Results of ablation studies on several variants of ConfCalib.
4.2.1. Main Results
The main results on each dataset are shown in Table 2. Our method shows better calibration performance compared with the competing methods. It achieves lower calibration errors on both field-wise and global metrics. Besides, our method also achieves higher AUC in all the tasks, indicating improvement in ranking ability in contrast to the uncalibrated model and other methods that may change the original order.
By contrast, the performance of scaling and binning-based comparative methods is unsatisfactory. These methods reduce global errors to some extent through global calibration, but the performance is not outstanding. Meanwhile, they are basically powerless against field-wise calibration errors.
Field-level calibration methods with neural networks such as NeuCalib and AdaCalib achieve relatively low calibration errors on the single target field among comparative methods. However, they are still inferior to our method, especially on multi-field metrics like Multi-field-RCE and global metrics like ECE, demonstrating the effectiveness of our multi-field joint calibration. Moreover, the results of the two neural methods show large variance, which means significantly unstable learning processes on small calibration datasets. On the contrary, our method only relies on posterior statistical information and maintains a stabler calibration effect.
It is noteworthy that results of the neural method AdaCalib on industrial data and AliExpress indicate that it may be unable to learn effectively in some cases, resulting in AUC even lower than the uncalibrated one. This phenomenon is likely caused by severe data sparsity. After sequential divisions of the feature value split and the binning operation, the number of total and positive samples within each subset is very limited. Thus, it is difficult for AdaCalib to learn the network parameters of corresponding parts of the piece-wise calibration curves. Our method works well in the same situation, benefiting from the awareness of confidence and the multi-field fusion approach designed to deal with sparse data.
4.2.2. Ablation Study
We report the results of several variants of our ConfCalib on Avazu CTR task in Fig. 5. Concretely, we respectively remove the confidence-aware calibration with Wilson interval, the nonlinear z-score mapping function (replaced by a single linear mapping), and the multi-field joint calibration from the original ConfCalib. The results show that all variants have varying degrees of increase in calibration errors and decrease in AUC, which can demonstrate the effectiveness of each key point.
The confidence-aware calibration partly preserves confidence in original predictions while calibrating based on observed posterior probabilities, which can avoid overfitting on limited validation samples in case the bias direction in the test dataset is opposite. In other words, it can effectively mitigate the impact of real-time data shifts during online serving. Although we cannot simulate frequent data shifts using a fixed test dataset, there is still an evident decline in AUC and an increase in Field-RCE after removing the confidence-aware calibration.
The elaborate z-score mapping function dynamically adjusts the calibration intensity according to the deviations between predicted and observed probabilities. For samples with large deviations, it introduces an upper bound to force the calibrated result back to a high enough confidence level. After replacing the mapping function with a single linear function $g(z)=\lambda z$, AUC and all calibration error metrics significantly deteriorate. The main reason is that the unbounded linear function chooses insufficient calibration strength for samples with large deviations.
The multi-field joint calibration simultaneously considers the biases across multiple fields and conducts joint calibration with optimized weights of each field. Compared to calibration on a single field, the joint calibration on multiple fields significantly improves both field-wise and global calibration errors, as well as AUC.
Table 3. Results of recalibration on comparative methods.
Method AUCnormal-$\uparrow$ Field-RCEnormal-$\downarrow$ Multi-field-RCEnormal-$\downarrow$ ECEnormal-$\downarrow$ MVCEnormal-$\downarrow$
GaussCalib
0.7545
0.1969
0.1470
0.0058
0.0041
+ConfCalib 0.7577 0.1417 0.1105 0.0053
0.0050
GammaCalib
0.7545
0.1984
0.1480
0.0058
0.0041
+ConfCalib 0.7577 0.1430 0.1113 0.0053
0.0050
NeuCalib
0.7554
0.1491
0.1091
0.0110
0.0042
@@ 4.2.2. Ablation Study
AdaCalib
0.7569
0.1650
0.1308
0.0099
0.0065
+ConfCalib
0.7569
0.1535 0.1069 0.0072 0.0024
4.3. Recalibration on Comparative Methods
Part of the comparative methods are learning-based, such as GaussCalib, GammaCalib, NeuCalib, and AdaCalib. It can be seen from the results in Table 2 that the performances of these methods are not satisfying for remaining high calibration errors or reducing AUC. Since our method does not require extra model training and has a low deployment cost, it can recalibrate the calibrated output of other methods for further improvement.
We experiment on Avazu CTR task. In detail, we first independently fit our method and the comparative methods to be recalibrated on the validation set. Then, the comparative methods make calibrated predictions for the test set. Finally, our method conducts recalibration on the calibrated outputs. Table 3 shows that recalibration using our ConfCalib achieves further improvement compared with the original results of other calibration methods. Profiting from the multi-field joint calibration, the recalibration significantly reduces Multi-field-RCE and the global calibration error MVCE. Besides, there is a large increase in AUC for scaling-based methods, including GaussCalib and GammaCalib, which strictly keep the order of the original predictions.
@@ 4.4. Robustness Analysis against Data Sparsity
As described in Section 3.2, one main advantage of ConfCalib is its robustness against data sparsity. Due to the limited time window for online data collection, the number of samples available to fit calibration models can be very few. Under such circumstances, methods based on neural networks can lead to inferior effects, although they may behave well if trained on sufficient samples.
Observing the results on Avazu in Table 2, we can find that the neural methods NeuCalib and AdaCalib both achieve lower calibration errors compared with the original predictions. Here, we reduce the number of validation samples in Avazu with various sample rates to simulate scenarios with varying degrees of data sparsity. The calibration performances w.r.t different sparsity degrees are shown in Fig. 6. As the number of samples for fitting the calibration model decreases, the performances of NeuCalib, AdaCalib, and ConfCalib all deteriorate in general. Fluctuations in the results of the two neural methods are intense, indicating that they are susceptible to data sparsity. However, our ConfCalib is little affected and exhibits strong robustness against data sparsity.
@@ 4.5. Analysis of Multi-field Fusion Weights
The multi-field joint calibration of ConfCalib obtains fusion weights for multiple target fields by minimizing the calibration error on single or multiple fields. Here, we select three target fields, ”ad_type”, ”consume_frequency”, and ”price” in our industrial dataset, and optimize Field-RCE of each field and Multi-field-RCE using grid search to acquire four groups of fusion weights separately, as shown in Fig. 7. Each column takes Field-RCE of the corresponding field as the optimization object. We can see that each field makes effects with different weights when optimizing Multi-field-RCE of all three fields. When optimizing Field-RCE on a single field, the weights of other fields are not zero, indicating that joint calibration on multiple fields can also reduce the single-field calibration error. The weight of ”ad_type” is generally the highest, suggesting that this field can better reflect global biases. Nevertheless, when Field-RCE on ”ad_type” is minimized, the weight of the other field ”consume_frequency” is the highest instead. This phenomenon indicates that calibration on a single field may not be sufficient, while calibration on other fields can provide effective assistance.
Figure 7. Fusion weights in multi-field joint calibration.
@@ 4.6. Hyperparameter Analysis
Our method has a single hyperparameter $\lambda$ located in the z-score mapping function to control the calibration strength manually. Its main effect is to select a relatively higher confidence level when observed samples are adequate, thus making the calibrated prediction closer to the observation. Here, we discuss the influence of different settings of $\lambda$. In detail, we set $\lambda\in\{0,0.2,0.4\dots,3.0\}$ and perform experiments on the industrial dataset. The results are shown in Fig. 8. Overall, the fluctuation of field-wise calibration errors like Field-RCE and Multi-field-RCE is within a small range as $\lambda$ changes. For global calibration error metrics, the change of ECE is relatively violent when $\lambda$ is large, while MVCE remains almost unchanged throughout. When $\lambda=0$, the awareness of confidence is removed, which means we disregard the model prediction and believe entirely in the observation. We achieve a local optimum at this point. As $\lambda$ increases, the calibration intensity gradually decreases, and the global optimal $\lambda$ is around 1.6, considering all the calibration errors.
@@ 4.7. Online A/B Test
We conduct one week’s online A/B test separately on two business scenarios to evaluate the effectiveness of ConfCalib in practical applications. In each scenario, we randomly split users into two groups, each with tens of millions of users everyday. One group receives recommendations by ConfCalib, while the other one gets recommendations by a well-crafted baseline. On an industrial advertising platform, ConfCalib achieves a relative improvement of 2.42% in CVR compared with the baseline. On the Top-grossing of a popular Android app store, ConfCalib obtains relative improvement of 32.6% and 49.1% in CTR and the revenue over the baseline. The data distribution shifts frequently on the Top Grossing scenario, and ConfCalib can better adapt this variability to achieve a significant improvement.
@@ 5. Conclusion
In this paper, we propose a confidence-aware multi-field model calibration method, ConfCalib, to address the common phenomenon of data sparsity in model calibration on advertising platforms. ConfCalib can dynamically adjust the calibration intensity based on the number of observed samples and posterior probabilities while considering the biases along multiple feature fields for joint calibration to capture various bias patterns and further alleviate the impact of data sparsity. We conduct sufficient comparative experiments on public and industrial datasets of different user action prediction tasks, and the results demonstrate that ConfCalib outperforms prior methods and exhibits strong robustness against data sparsity. ConfCalib is currently running on an industrial advertising platform and a popular Android app store, and has achieved continuous online revenue improvement.

### v2

@@ 4.1.4. Evaluation Metrics
[…]
@@
Figure 5. Results of the ablation study on several variants of ConfCalib.
4.2. Results
4.2.1. Main Results
The main results on each dataset are shown in Table 2. Our method shows better calibration performance compared with baselines. It achieves lower calibration errors on both field-wise and global metrics compared to the uncalibrated model and almost all the baselines. Besides, our method also achieves higher AUC, indicating an improvement in ranking ability. Scaling and binning-based baselines show poor performances. They reduce global errors to some extent through global calibration while powerless against field-wise calibration errors. Field-level calibration methods such as NeuCalib and AdaCalib achieve low calibration errors on the single target field among baselines. However, they are still inferior to our method, especially on multi-field metrics like Multi-Field-RCE and global metrics like ECE, demonstrating the effectiveness of our multi-field joint calibration. The two neural methods show large variance, which means unstable learning on sparse data. On the contrary, our method only relies on posterior statistical information and maintains a stable calibration effect. Moreover, AdaCalib obtains a lower AUC than the uncalibrated model on our industrial dataset, indicating that complex neural calibration methods may be counterproductive under data sparsity.
4.2.2. Ablation Study
We report the results of several variants of our ConfCalib on Avazu CTR task in Fig. 5. Concretely, we respectively remove the confidence-aware calibration (i.e., turn into a field-wise naive calibration), the nonlinear z-score mapping function (replaced by a single linear mapping), and the multi-field joint calibration from the original ConfCalib. The removal of our confidence-aware calibration results in increase in field-wise calibration errors and decrease in AUC. After replacing the mapping function with a single linear function $g(z)=\lambda z$, AUC and all calibration error metrics deteriorate. This validates that the bounded mapping can provide proper calibration intensity compared with simple scaling. Finally, removing the joint calibration on multiple fields significantly increases both field-wise and global calibration errors and causes a large decline in AUC.
Table 3. Results of recalibration on learning-based baselines.
@@ 4.2.2. Ablation Study
GaussCalib 0.7545 0.1969 0.1470 0.0058 0.0041
+ConfCalib 0.7577 0.1417 0.1105 0.0053 0.0050
GammaCalib 0.7545 0.1984 0.1480 0.0058 0.0041
+ConfCalib 0.7577 0.1430 0.1113 0.0053 0.0050
NeuCalib 0.7554 0.1491 0.1091 0.0110 0.0042
@@
AdaCalib 0.7569 0.1650 0.1308 0.0099 0.0065
+ConfCalib 0.7569 0.1535 0.1069 0.0072 0.0024
4.3. Recalibration on Learning-Based Baselines
Part of the baselines are learning-based, such as GaussCalib, GammaCalib, NeuCalib, and AdaCalib. Since our ConfCalib does not require extra model training and has a low deployment cost, it can recalibrate the calibrated output of other methods for further improvement. We experiment on Avazu CTR task. First, we fit our method and baselines to be recalibrated separately on the validation set. Then, baselines make calibrated predictions for the test set. Finally, our method conducts recalibration on the calibrated outputs. Table 3 shows that recalibration achieves further improvement compared with the original results of other methods. Profiting from multi-field joint calibration, recalibration using ConfCalib significantly reduces Multi-Field-RCE and the global calibration error MVCE. Besides, recalibration on methods that strictly keep the order of original predictions, such as GaussCalib and GammaCalib, can provide an improvement in AUC as well.
@@ 4.4. Robustness Analysis against Data Sparsity
One primary advantage of ConfCalib is its robustness against data sparsity. Due to the limited time window for online data collection, samples available to fit calibration models can be very few. Under such circumstances, methods based on neural networks may be ineffective or even counterproductive. Here, we reduce the validation dataset used for calibration in Avazu with various sample rates to simulate scenarios with varying degrees of data sparsity. The calibration performances w.r.t different sparsity degrees are shown in Fig. 6. As the number of samples for fitting the calibration model decreases, the performances of NeuCalib, AdaCalib get worse evidently, while our ConfCalib is little affected and exhibits strong robustness against data sparsity.
Figure 7. Fusion weights in multi-field joint calibration.
@@ 4.5. Analysis of Multi-field Fusion Weights
The multi-field joint calibration of ConfCalib obtains fusion weights for multiple target fields by minimizing the calibration error on single or multiple fields. Here, we select three target fields, ”ad_type”, ”consume_frequency”, and ”price” in our industrial dataset, and optimize Field-RCE of each field and Multi-Field-RCE using grid search to acquire four groups of fusion weights separately, as shown in Fig. 7. Each column takes Field-RCE of the corresponding field as the optimization object. We can see that each field makes effects with different weights when optimizing Multi-Field-RCE of all three fields. When optimizing Field-RCE on a single field, the weights of other fields are not zero, indicating that joint calibration on multiple fields can also reduce the single-field calibration error. The weight of ”ad_type” is generally the highest, suggesting that this field can better reflect global biases. Nevertheless, when Field-RCE on ”ad_type” is minimized, the weight of the other field ”consume_frequency” is the highest instead. This indicates that calibration on a single field is insufficient, and multi-field calibration can provide effective assistance.
@@ 4.6. Hyperparameter Analysis
Our method has a single hyperparameter $\lambda$ in the z-score mapping function Eq. (2) to control the calibration strength manually. Its main effect is to select a relatively higher confidence level when observed samples are adequate, thus making the calibrated prediction closer to the observation. Here, we discuss the influence of different settings of $\lambda$. We set $\lambda=0,0.2,0.4\dots,3.0$ and experiment on the industrial dataset. Results are shown in Fig. 8. Overall, the variation of field-wise calibration errors like Field-RCE and Multi-Field-RCE is within a small range as $\lambda$ changes. For global calibration error metrics, the change of ECE is relatively violent when $\lambda$ is significant, while MVCE remains almost the same. When $\lambda=0$, the awareness of confidence is removed, which means we disregard the model prediction and believe entirely in the observation. We achieve a local optimum at this point. As $\lambda$ increases, the calibration intensity gradually decreases, and the global optimal $\lambda$ is around 1.6.
@@ 4.7. Online A/B Test
We conduct one week’s online A/B test separately on two industrial scenarios to evaluate ConfCalib’s effectiveness in practical applications. In each scenario, we randomly split users into two groups, each with tens of millions of users daily. One group receives recommendations from ConfCalib, while the other gets recommendations from a well-crafted baseline. On Huawei’s online advertising platform, ConfCalib achieves a relative improvement of 2.42% in CVR compared with the baseline. On Top-grossing of Huawei AppGallery, ConfCalib obtains relative improvement of 32.6% and 49.1% in CTR and revenue. Top-grossing is a small-scale scenario where data distribution shifts frequently, and ConfCalib can better adapt this variability to achieve a significant improvement.
@@ 5. Conclusion
In this paper, we propose a confidence-aware multi-field model calibration method, ConfCalib, to address the common phenomenon of data sparsity in model calibration for advertising recommendation. ConfCalib dynamically adjusts calibration intensity based on the confidence level obtained from the observed data distribution. Moreover, it considers biases along multiple feature fields and introduces joint calibration to capture various bias patterns and further alleviate the impact of data sparsity. We conduct sufficient offline and online experiments. Results demonstrate that ConfCalib outperforms prior methods and exhibits strong robustness against data sparsity. ConfCalib is currently running on Huawei’s advertising platform and AppGallery and has achieved continuous online revenue improvement.